### v1

# Skin Effect Induced Anomalous Dynamics from Charge-Fluctuating Initial States

Abstract
Non-equilibrium dynamics in non-Hermitian systems has attracted significant interest, particularly due to the skin effect and its associated anomalous phenomena. Previous studies have primarily focused on initial states with a definite particle number. Here, we present a systematic study of non-reciprocal quench dynamics in the pairing states with indefinite particle number. Our study uncovers a range of novel behaviors. Firstly, we demonstrate a universal tendency towards half-filling of particle density at late times. At early times for certain initial states, we observe a chiral wavefront in both particle number distribution and charge inflow, associated with a sharp decrease in particle number. Furthermore, we find that non-Hermiticity could enhance the growth of entanglement in the initial stages of evolution. In the intermediate time regime, the characteristic skin effect leads to particle accumulation on one side, leading to a pronounced reduction in entanglement entropy. Moreover, our results reveal the presence of the quantum Mpemba effect during the restoration of U(1) symmetry. Our findings open new avenues for exploring exotic dynamic phenomena in quantum many-body systems arising from the interplay of symmetry breaking and non-Hermiticity.
Introduction.—The study of non-equilibrium dynamics, especially in open systems, has recently emerged as a major research frontier, revealing intriguing phenomena beyond the scope of conventional equilibrium descriptions Diehl et al. (2008, 2011); Barontini et al. (2013); Dast et al. (2014); Huang et al. (2018); Li et al. (2019); Jian et al. (2020); Lapp et al. (2019); Fuji and Ashida (2020); Choi et al. (2020); Huang and Liu (2020); Takasu et al. (2020); Gullans and Huse (2020); Bao et al. (2020); Lu and Grover (2021); Sang and Hsieh (2021); Turkeshi et al. (2021); Bätge et al. (2021); Wald et al. (2021); Levy et al. (2021); Li et al. (2021); Turkeshi and Schiró (2023); Kawabata et al. (2023); Moreira et al. (2024); Paz et al. (2024); Lessa et al. (2025); Despres (2025); Gao and Muttalib (2025); Glatthard (2025); Capecelatro et al. (2025). Non-Hermitian Hamiltonians offer an effective theoretical framework for describing open quantum systems Ashida et al. (2020); Prosen (2008); Lieu et al. (2020); Prosen (2010); Zhou et al. (2020); Zhou and Yu (2022); Ren et al. (2022a). Such non-Hermiticity naturally emerges when a system couples to an environment, introducing dissipation, gain, or non-reciprocal interactions. Moreover, non-Hermitian Hamiltonians provide a conceptually valuable description for quasiparticles possessing finite lifetimes due to mechanisms like electron-phonon coupling or disorder scatteringShen and Fu (2018); Nagai et al. (2020); Huang et al. (2025). The investigation of non-Hermitian systems has unveiled a wealth of exotic phenomena lacking counterparts in Hermitian systems Rudner and Levitov (2009); Hu and Hughes (2011); Esaki et al. (2011); Heiss (2012); Leykam et al. (2017); Couvreur et al. (2017); Nakagawa et al. (2018); Gong et al. (2018); McDonald et al. (2018); Xiong (2018); El-Ganainy et al. (2018); Martinez Alvarez et al. (2018); Kawabata et al. (2019a); Liu et al. (2019); Lee et al. (2019a); Yamamoto et al. (2019); Pan et al. (2020); Zhai et al. (2020); Iskin (2021); Lv et al. (2022); Kawabata et al. (2021); Zhai et al. (2022); Liang et al. (2022); Lee and Thomale (2019); Lee et al. (2019b); Chang et al. (2020); Li et al. (2020); Bergholtz et al. (2021); Xi et al. (2021); Ding et al. (2022); Kozii and Fu (2024); Guo et al. (2023); Shen et al. (2023); Okuma and Sato (2023); Mao et al. (2024); Zhong et al. (2024); Guo et al. (2020); Yang et al. (2020); Sun et al. (2020); Hu and Zhao (2021); Li and Hu (2023); Zheng et al. (2024); Ding et al. (2023); Chen et al. (2024); Gao et al. (2022); Chen et al. (2022); Zhang et al. (2023a); Wang et al. (2022); Zhang et al. (2022a); Yang et al. (2024); Hu et al. (2023); Jing et al. (2024); Yu et al. (2024); Murshed and Roy (2024); Juričić and Roy (2024); Tang et al. (2024); Liu and Batchelor (2025) such as the celebrated non-Hermitian skin effect (NHSE) with macroscopic accumulation of eigenstates localized at the system’s edges Yao and Wang (2018); Yokomizo and Murakami (2019); Yao et al. (2018); Kawabata et al. (2019b); Yokomizo and Murakami (2019); Okuma et al. (2020); Zhang et al. (2020); Song et al. (2019); Helbig et al. (2020); Kawabata et al. (2020); Zou et al. (2021); Zhou et al. (2022); Longhi (2021); Kim and Park (2021); Yokomizo and Murakami (2021); Mao et al. (2021); Xiao et al. (2020); Zhang et al. (2022b); Guo et al. (2022); Feng et al. (2025); Sun et al. (2021); Li et al. (2023a); Yoshida et al. (2024); Huang et al. (2024); Wang et al. (2024a); Zhang et al. (2024); Hu (2025); Sun et al. (2025).
While significant progress has been made in understanding non-Hermitian dynamics, much of the focus has been on systems with particle-number conservation. However, many physical processes particularly those involving Cooper pairing and superconductivity, inherently lead to states that lack a definite particle number, breaking U(1) symmetry. The non-equilibrium dynamics of such U(1)-breaking states under non-Hermitian evolution, and the interplay between particle number fluctuations and unique non-Hermitian features like the NHSE, remain largely unexplored. This regime presents fundamental challenges and opportunities to discover novel phenomena arising from the confluence of non-unitary dynamics, non-reciprocity, and symmetry breaking.
Probing the non-equilibrium evolution of quantum many-body states requires analyzing multiple facets of their properties. Fundamental quantities like particle density provide essential information about the spatial distribution and transport of particles, crucial for understanding phenomena like NHSE. Beyond density, quantum entanglement encapsulates the unique non-classical correlations fundamental to quantum systems Eisert et al. (2010); Calabrese and Cardy (2004); Alba and Calabrese (2017); Ashida and Ueda (2018); Skinner et al. (2019); Herviou et al. (2019); Chen et al. (2020); Bácsi and Dóra (2021); Orito and Imura (2022); Dóra et al. (2022); Turkeshi and Schiró (2023); Kawabata et al. (2023); Orito and Imura (2023); Li et al. (2023b, 2024). Monitoring its evolution, often quantified as entanglement entropy (EE), provides a valuable dynamic window into a system’s properties Cao et al. (2019); Fuji and Ashida (2020); Alberton et al. (2021); Turkeshi et al. (2021); Kells et al. (2023); Wang et al. (2024b). Furthermore, to specifically track the restoration of broken U(1) symmetry, entanglement asymmetry (EA) has been introduced as a sensitive probe derived from entanglement, capable of detecting subtle changes in the symmetry properties of the state  Ares et al. (2023a). Analyzing EA dynamics in Hermitian systems has revealed intriguing symmetry restoration phenomena, including the quantum Mpemba effect (QME) Nava and Fabrizio (2019); Carollo et al. (2021); Zhang et al. (2025); Moroder et al. (2024); Bettmann and Goold (2025); Chatterjee et al. (2023, 2024); Aharony Shapira et al. (2024); Manikandan (2021); Kochsiek et al. (2022); Ivander et al. (2023); Zhou et al. (2023); Wang and Wang (2024); Liu et al. (2024a); Longhi (2024); Nava and Egger (2024); Boubakour et al. (2025); Foligno et al. (2025); Zatsarynna et al. (2025); Longhi (2025); Dong et al. (2025); Graf et al. (2025); Van Vu and Hayakawa (2025); Teza et al. (2025); Joshi et al. (2024); Chang et al. (2024); Caceffo et al. (2024); Strachan et al. (2024), where states initially exhibiting stronger symmetry breaking can paradoxically restore symmetry faster Ares et al. (2023a, b); Murciano et al. (2024); Chalas et al. (2024); Yamashika et al. (2024a); Rylands et al. (2024a); Yamashika et al. (2024b); Rylands et al. (2024b); Liu et al. (2024b, c); Yu et al. (2025); Giulio et al. (2025); Klobas et al. (2025); Ares et al. (2025); Ferro et al. (2024); Klobas (2024). While these quantities have been extensively studied in various contexts, their interplay and behavior under non-Hermitian evolution, particularly for initial states lacking particle-number conservation, remain elusive. This motivates fundamental questions: How are particle density profiles, entanglement growth, and U(1) symmetry restoration affected by non-Hermitian dynamics and the NHSE in states without definite particle number?
To address these essential questions, we investigate the non-equilibrium dynamics of a superconducting pairing state under the time evolution of a typical non-Hermitian Hamiltonian, termed the Hatano-Nelson (HN) model Hatano and Nelson (1997), which involves non-reciprocal hopping of fermions exhibiting the skin effect. For the time evolution of the pairing state under HN model, the total particle number is not conserved. Here, we demonstrate a universal late time convergence to half-filling, arising from the skin effect, regardless of the initial particle number. Furthermore, we uncover a range of novel phenomena in the short-time dynamics arising from the interplay of U(1) symmetry breaking and non-Hermiticity. (1) We observe a chiral wavefront in the particle density dynamics, which is also manifested in the behavior of particle inflow due to non-Hermiticity for certain initial states. (2) The rate of entanglement growth is enhanced at early times while greatly suppressed at late times by the non-Hermiticity. (3) In the process of U(1) symmetry restoration, QME persists with the introduction of non-Hermiticity. Crucially, these universal results are robust across a wide range of parameters within the model, holding significant promise for experimental verification.
Model and Method.—We employ the initial state with charge fluctuation by considering the ground state of the one-dimensional fermionic Hamiltonian involving Cooper pairing:
$\hat{H}_{\text{ini}}=-\sum_{j}(J\hat{c}_{j}^{\dagger}\hat{c}_{j+1}+\Delta\hat{c}_{j}^{\dagger}\hat{c}_{j+1}^{\dagger}+\text{h.c.})-\mu\sum_{j}\hat{c}_{j}^{\dagger}\hat{c}_{j}.$ (1)
where $\hat{c}_{j}^{\dagger}$ ($\hat{c}_{j}$) represent the creation (annihilation) operators, $J$ is the amplitude of nearest-neighbor (NN) hopping, $\mu$ is chemical potential, and $\Delta$ denotes the pairing amplitude. The initial state is thus a superconducting BCS pairing state breaking charge U(1) symmetry. Through Jordan-Wigner transformation (see the section I of the supplemrental material (SM) Ref for the details), the fermionic Hamiltonian in Eq. (1) is straightforwardly transformed to a spin-1/2 chain as follows:
$\hat{H}_{\text{ini}}=-$ $\frac{J+\Delta}{2}\sum_{j}\hat{\sigma}_{j}^{x}\hat{\sigma}_{j+1}^{x}-\frac{J-\Delta}{2}\sum_{j}\hat{\sigma}_{j}^{y}\hat{\sigma}_{j+1}^{y}$ (2)
-- $\frac{\mu}{2}\sum_{j}\hat{\sigma}_{j}^{z}.$
$\Delta$ breaks U(1) spin rotational symmetry around $z$-axis. We implement anti-periodic boundary condition (APBC) of Eq. (1), corresponding the periodic boundary condition (PBC) of spin-1/2 chain in Eq. (2) under Jordan-Wigner transformation for even number of lattice sites. The ground state of Eq. (1) is conveniently expressed in the spin representation as a superposition of tilted ferromagnetic states Ares et al. (2023a):
$|\psi(0)\rangle=$ $\frac{1}{\sqrt{2}}(|F,\theta\rangle+|F,-\theta\rangle),$ (3)
[…]
where $\theta$ represents the tilted angle characterizing the strength of U(1) symmetry breaking. The relations between $\theta$, $\Delta$ and $\mu$ satisfy Ares et al. (2023a, b)
$\text{cos}^{2}\theta=\frac{1-\Delta}{1+\Delta},\Delta^{2}+\mu^{2}=1.$ (4)
To investigate the non-Hermitian dynamics, we employ the HN model Hatano and Nelson (1997) whose Hamiltonian is:
$\hat{H}_{\text{evo}}=-\sum_{j}(t_{L}\hat{c}_{j}^{\dagger}\hat{c}_{j+1}+t_{R}\hat{c}_{j+1}^{\dagger}\hat{c}_{j}),$ (5)
where $t_{L}=J+\gamma$, $t_{R}=J-\gamma$ are the non-reciprocal hopping coefficients. Crucially, the non-reciprocity ($\gamma\neq 0$) inherent in these hopping coefficients is responsible for the skin effect under open boundary conditions (OBC).
For non-unitary time evolution, the normalized state $|\psi(t)\rangle$ at time $t$ is given by:
$|\psi(t)\rangle=\frac{e^{-i\hat{H}_{\text{evo}}t}|\psi(0)\rangle}{\sqrt{\langle\psi(0)|e^{i\hat{H}_{\text{evo}}^{\dagger}t}e^{-i\hat{H}_{\text{evo}}t}|\psi(0)\rangle}}.$ (6)
For numerical calculations, we use the quantum software TensorCircuit-NG *[][.https://github.com/tensorcircuit/tensorcircuit-ng.]Zhang2022_z. The energy unit is set as $J=1$, system size $L=64$ and OBC are employed. Without loss of generality, we focus on $\gamma\geq 0$, which results in the skin effect where particles accumulate at the left boundary. It is important to note that the combination of non-reciprocity and the large system size can introduce significant numerical errors in the direct calculation of $e^{-i\hat{H}_{\text{evo}}t}$ Feng et al. (2025). To maintain numerical accuracy, we utilize a small time-step decomposition of the evolution and perform calculations with a high precision of 250 digits (See the details in section II of the SM Ref ).
Figure 1: (a) Evolution of the total particle number starting from various initial tilted ferromagnetic states. (b) Dynamics of local particle density $n_{j}$. Characteristic timescales $\tau_{1}$ and $\tau_{2}$ are marked by horizontal dashed lines. Dynamics of (c) current density $I_{j}$, and (d) particle inflow rate $\sigma_{j}$. In (c), positive current indicates flow from left to right, and negative current from right to left. In (d), positive inflow represents particle gain, and negative inflow represents particle loss. For (b)-(d), the initial state is the tilted ferromagnetic state with $\theta=\pi/6$. The system size is $L=64$ and the non-reciprocity strength is $\gamma=0.8$ for all figures.
Figure 2: Entanglement entropy under open boundary conditions. (a), (b) and (c) correspond to $\gamma=0.0$, 0.2 and 0.6, respectively, where the legends represent different tilted angles. The traced subsystem $B$ with length $l=6$ is located at the left end of the system. The total system size is $L=64$.
Universal late-time half-filling convergence. Different from the initial state with particle-number conservation, a prominent feature of the U(1) symmetry breaking state under time evolution of HN model is that the particle number is not conserved. We first examine the dynamics of the total particle number, defined by
$\overline{N}(t)=\sum_{j}\langle\psi(t)|\hat{n}_{j}|\psi(t)\rangle,$ (7)
where $\hat{n}_{j}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the on-site particle density. The results of $\overline{N}(t)$ for the tilted ferromagnetic state with different $\theta$ and fixed $\gamma=0.8$, shown in Fig. 1(a), decrease with time. Most notably, we uncover a universal long-time behavior: regardless of their initial particle number, all tilted ferromagnetic states ($\theta>0$) converge to a total particle number close to the half filling ($L/2$) under long-time evolution. We provide further justification for this observation, including quantitative analysis and a rigorous proof in section III of the SM Ref . Specifically, we rigorously demonstrate that in the limit $\gamma\rightarrow 1$ for this HN model, the total particle number asymptotically approaches $L/2$ as $t\to\infty$. In contrast, the ferromagnetic state ($\theta=0$), as an eigenstate of the on-site particle density operator $\hat{n}_{j}$, remains fully occupied during the evolution. In terms of systems with PBC, the total particle number for a general initial state also tends towards $L/2$ in the steady state.
Density dynamics.— To further explore the dynamical characteristics reflected in the particle number, we now examine the time evolution of the on-site particle density $\langle\hat{n}_{j}\rangle$, exemplified in Fig. 1(b) for $\theta=\pi/6$ Ref . While the initial state exhibits a uniform particle density, distinct behaviors emerge at different timescales. At long times, there is a clear spatial separation of particles due to the skin effect. The particle density becomes concentrated on the left side of the system, leaving the opposite side sparsely populated. Intriguingly, short-time dynamics reveal a contrasting exotic behavior. Instead of accumulation, we observe an initial reduction of particle density near the left edge. This sharp depletion is linked to a chiral wavefront propagating with twice the maximum group velocity of quasiparticles. Specifically, particle density begins to decrease as the wavefront reaches a given site. Here, the group velocity is defined as $v_{\text{g}}=\frac{\mathrm{d}}{\mathrm{d}k}\text{Re}[E(k)]$, where $E(k)=-2J\text{cos}k-2i\gamma\text{sin}k$ is the energy spectrum of the HN model under PBC Modak et al. (2020); Orito and Imura (2023); Cevolani et al. (2018).
With increasing time, two characteristic timescales, $\tau_{1}$ and $\tau_{2}$, become evident. The first $\tau_{1}$ corresponds to the arrival time of the chiral wavefront at the right boundary, at which point the total particle number approaches $L/2$ for large $\gamma$. Around this time, the particle density begins to change from both boundaries inwards: increasing from the left edge and decreasing from the right edge towards the center. Upon reaching the second timescale, $\tau_{2}$, the density distribution stabilizes into a steady state, with only minor oscillations remaining in the central region. This steady state distribution clearly demonstrates the skin effect, with the left half of the system almost fully occupied and the right half nearly empty. In stark contrast, an initial tilted antiferromagnetic (AFM) state lacks this chiral wavefront; its dynamics immediately resemble the post-$\tau_{2}$ behavior of the tilted ferromagnetic state, likely because the AFM state starts at half-filling. Detailed results are presented in section V of the SM Ref .
The time-evolution of particle density is determined by the continuity equation as:
$\frac{\partial n_{j}}{\partial t}+(I_{j}-I_{j-1})=\sigma_{j},$ (8)
where $I_{j}=iJ(\hat{c}^{\dagger}_{j+1}\hat{c}_{j}-\hat{c}^{\dagger}_{j}\hat{c}_{j+1})$ represents the current flowing out of site $j$ and into site $j+1$, and $\sigma_{j}$ denotes the inflow of particles into site $j$ from the external environment. In Hermitian systems without particle dissipation, $\sigma_{j}$ vanishes, and Eq. (8) reduces to the standard continuity equation. To gain further insight into the particle dynamics, particularly the influence of current density and particle inflow, we show the dynamical results for current density and particle inflow in Fig. 1(c) and (d), respectively.
Similar to the particle density, characteristic timescales $\tau_{1}$ and $\tau_{2}$ are also evident in the dynamics of both current density and particle inflow. In the short-time regime ($\tau<\tau_{1}$), a chiral wavefront emerges in both the current density and particle inflow profiles. The characteristic time $\tau_{1}$ corresponds to the moment when this chiral wavefront reaches the right boundary of the system. Notably, the wavefront in the particle inflow in Fig. 1(d) represents a significant particle sink, indicating a substantial loss of particle number at given space-time location. The wavefront of current in Fig. 1(c) appears as a sharp spatial gradient of current density. Although the negative current gradient implies particle accumulation according to the continuity equation Eq. (8), the observed net effect at the wavefront is particle depletion. Thus, the particle number reduction associated with the wavefront in Fig. 1(b) arises from the effect of particle inflow acting as a propagating sink, which dominates over the particle-accumulating effect of the current induced by non-reciprocity. In the intermediate stage ($\tau_{1}<\tau<\tau_{2}$), the rapid increase in particle number at the left edge and decrease at the right edge are also driven by the particle inflow, as visualized in Fig. 1(d). For late times ($\tau>\tau_{2}$), both the current and particle inflow diminish, exhibiting only small, time-decaying oscillations. Results for other values of $\gamma$ show qualitatively consistent behaviors, as included in section V of the SM Ref . Moreover, our analysis of the density dynamics under APBC, detailed in the the section VI of the SM Ref , yields qualitatively distinct conclusions, further demonstrating that the anomalous phenomena observed under OBC are attributed to the skin effect.
Entanglement dynamics.—We now turn to the dynamics of bipartite entanglement, where we employ the bipartite von Neumann EE, defined as:
$S_{\text{vonEE}}=-\text{Tr}_{A}[\rho_{A}\text{ln}\rho_{A}].$ (9)
Here, the system is bipartitioned into subsystems $A$ and $B$, and $\rho_{A}=\text{Tr}_{B}[\rho]$ is the reduced density matrix of subsystem $A$. For the Hermitian case (Fig. 2(a)), the EE for different initial tilted states initially increases and subsequently saturates to a steady-state value. In contrast, the non-Hermitian dynamics, shown in Fig. 2(b) and (c), exhibit more complex phenomena. Similar to the particle density dynamics, the entanglement evolution displays distinct behaviors in different temporal regimes. In the stage $\tau<\tau_{1}$, EE increases with evolution time and subsequently saturates, qualitatively consistent with the conventional behavior of entanglement dynamics in Hermitian systems. This trend is qualitatively captured by the quasiparticle picture, as demonstrated in section IV of the SM Ref . An intriguing observation, illustrated in Fig. 2(b)-(c), is that non-Hermiticity can enhance both the growth rate and the maximum value of EE during time evolution for specific initial states and small $\gamma$.
In contrast to Hermitian systems, the EE exhibits a rapid decrease in the temporal regime $t>\tau_{1}$, a behavior fundamentally different from the Hermitian case. This is directly attributed to the particle density dynamics driven by the skin effect. As shown in Fig. 1(b), the skin effect leads to particle accumulation on the left side, with the occupied region expanding over time. Consequently, near the boundary between subsystems A and B, the system increasingly resembles a product state, which inherently suppresses entanglement. Thus, the observed EE reduction is a direct consequence of the particle number distribution arising from the skin effect.
Given the skin effect, we anticipate that the EE depends on the bipartition subsystem’s location. To investigate this, we calculate the EE for different subsystem positions, with detailed results presented in section V of the SM Ref . When the small subsystem is located centrally, the qualitative behavior of EE dynamics remains similar to the cases where the subsystem is at the left or right boundary. However, the reduction in EE after $t>\tau_{1}$ is significantly less pronounced compared to the results in Fig. 2(b) and (c). This milder reduction can be naturally understood through the particle number distribution induced by the skin effect, as the particle density fluctuation is stronger at the middle region.
Figure 3: Entanglement asymmetry under the open boundary conditions. (a), (b) and (c) correspond to $\gamma=0.0$, $0.2$ and $0.6$, respectively, where the legends represent different tilted angles. The traced subsystem $B$ with length $l=12$ is located at the left end of the system. The total system size is $L=64$.
[…]
$\Delta S^{(n)}_{A}=S_{n}(\rho_{A,N})-S_{n}(\rho_{A}),$ (10)
[…]
Recent research has revealed a novel and universal QME in dynamic symmetry restoration, where systems with greater initial symmetry breaking recover symmetry faster Ares et al. (2023a); Murciano et al. (2024); Caceffo et al. (2024); Giulio et al. (2025); Ares et al. (2023b); Chalas et al. (2024); Yamashika et al. (2024a); Rylands et al. (2024a); Yamashika et al. (2024b); Rylands et al. (2024b); Liu et al. (2024b, c); Yu et al. (2025); Chang et al. (2024). A key question is whether symmetry restoration and the QME can emerge in non-Hermitian systems. Fig. 3 presents the second-order Rényi EA for varying non-Hermiticity strengths ($\gamma$). Consistent with prior work Ares et al. (2023a), Fig. 3(a) shows a clear curve crossing for different tilted angles ($\theta$) in the Hermitian case, confirming the QME. Notably, a pronounced QME also exists in the non-Hermitian scenario, as evidenced by curve crossings in Fig. 3(b) and (c). Intriguingly, this QME appears enhanced compared to the Hermitian case, with curve crossings occurring earlier in time. This enhancement is accompanied by non-monotonic behavior in the early-time EA evolution for some initial states, where EA initially increases before decreasing towards symmetry restoration. At late times, the EA results show a restoration of U(1) symmetry, mirroring the behavior observed in Hermitian systems.
Conclusions and discussions.—In this Letter, we investigate the quench dynamics of the pairing state with indefinite particle number under the time evolution of HN model. For the first time, we demonstrate the emergence of universal half-filling behavior in the late-time limit of the HN model. Beyond this, we unveil a rich array of novel phenomena stemming from the interplay between U(1) symmetry breaking and non-Hermiticity. Specifically, we observe a chiral wavefront in the particle density dynamics, non-Hermiticity enhanced entanglement growth and QME. Our findings open new routes for understanding quantum many-body dynamics in the presence of both symmetry breaking and dissipation.
[…]
Acknowledgments.—-SG and ZXL is supported by the National Natural Science Foundation of China under the Grant No.12347107 and Grant No.12474146. SXZ is supported by a start-up grant at IOP-CAS.